# Variational Distillation of Diffusion Policies into Mixture of Experts

### v1

@@ Abstract
This work introduces Variational Diffusion Distillation (VDD), a novel method that distills denoising diffusion policies into Mixtures of Experts (MoE) through variational inference. Diffusion Models are the current state-of-the-art in generative modeling due to their exceptional ability to accurately learn and represent complex, multi-modal distributions. This ability allows Diffusion Models to replicate the inherent diversity in human behavior, making them the preferred models in behavior learning such as Learning from Human Demonstrations (LfD). However, diffusion models come with some drawbacks, including the intractability of likelihoods and long inference times due to their iterative sampling process. The inference times, in particular, pose a significant challenge to real-time applications such as robot control. In contrast, MoEs effectively address the aforementioned issues while retaining the ability to represent complex distributions but are notoriously difficult to train. VDD is the first method that distills pre-trained diffusion models into MoE models, and hence, combines the expressiveness of Diffusion Models with the benefits of Mixture Models. Specifically, VDD leverages a decompositional upper bound of the variational objective that allows the training of each expert separately, resulting in a robust optimization scheme for MoEs. VDD demonstrates across nine complex behavior learning tasks, that it is able to: i) accurately distill complex distributions learned by the diffusion model, ii) outperform existing state-of-the-art distillation methods, and iii) surpass conventional methods for training MoE.
@@ 1 Introduction
To obtain the benefits of both models, i.e., learning highly accurate generative models with diffusion and obtaining simple, tractable models using a mixture of experts, this work introduces Variational Diffusion Distillation (VDD), a novel method that distills diffusion models to MoEs. Starting from the variational inference objective, we derive a lower bound that decomposes the objective into separate per-expert objectives, resulting in a robust optimization scheme. Each per-expert objective elegantly leverages the gradient of the pre-trained score function such that the MoE benefits from the diffusion model’s properties. The resulting MoE policy performs on par with the diffusion model and covers the same modes, while being interpretable, faster during inference, and has a tractable likelihood. This final policy is readily available to the user for post hoc analysis or fast fine-tuning for more specific situations. A high-level architecture of the VDD model and its relation to the diffusion policy is shown in Fig. 1. VDD is thoroughly evaluated on nine complex behavior-learning tasks that demonstrate the aforementioned properties. As an additional insight, this paper observed that one-step continuous diffusion models already perform well, a finding not discussed in prior work.
@@ 2 Related Works
Knowledge Distillation from Diffusion Models Knowledge distillation from diffusion models has been researched in various research areas. For instance, in text-to-3D modeling, training a NeRF-based text-to-3D model without any 3D data by mapping the 3D scene to a 2D image and leveraging a text-to-2D diffusion is proposed [7]. The work proposes minimizing the Score Distillation Sampling (SDS) loss that is inspired by probability density distillation [38] and incentivizes the 3D-model to be updated into higher density regions as indicated by the score function of the diffusion model. To overcome drawbacks such as over-smoothing and low-diversity problems when using the SDS loss, variational score distillation (VSD) treats the 3D scene as a random variable and optimizes a distribution over these scenes such that the projected 2D image aligns with the 2D diffusion model [8]. In a similar context, the work in [39] proposes distilling a trained diffusion model into another diffusion model while progressively reducing the number of steps. However, even though the number of diffusion steps is drastically reduced, a complete distillation, i.e. one-step inference as for VDD is not provided. Additionally, the resulting model suffers from the same drawbacks of diffusion models such as intractable likelihoods. In contrast, in consistency distillation (CD), diffusion models are distilled to consistency models (CM) [40, 41, 42, 43] such that data generation is possible in one step from noise to data. However, one-step data generation typically results in lower sample quality, requiring a trade-off between iterative and single-step generation based on the desired outcome. As CMs, VDD performs one-step data generation but distills the pre-trained diffusion model to an MoE which has a tractable likelihood and is efficient in inference time. The experimental evaluations show the advantages of VDD over CMs. Score Regularized Policy Optimization (SRPO) [44] also leverages a diffusion behavior policy to regularize the offline RL-based objective. However, in contrast to SRPO, VDD learns an MoE policy instead of a uni-modal Gaussian policy and explicitly distills a diffusion model instead of using it as guidance during optimization. Furthermore, VDD trains MoEs policies in imitation learning instead of reward-labeled data as in offline RL.
@@ 3 Preliminaries
where $q^{\xi}(z|\mathbf{s})$ and $q^{\nu_{z}}(\mathbf{a}|\mathbf{s},z)$ are referred to as gating and experts respectively. $\xi$ and $\nu_{z}$ denote the gating and expert parameters and thus $\phi=\xi\cup\{\nu_{z}\}_{z}$ The gating is responsible for soft-partitioning the state space into sub-regions where the corresponding experts approximate the target density. To sample actions, that is, $\mathbf{a}^{\prime}\sim q^{\phi}(\cdot|\mathbf{s}^{\prime})$ for some state $\mathbf{s}^{\prime}$, we first sample a component index from the gating, i.e., $z^{\prime}\sim q^{\xi}(\cdot|\mathbf{s}^{\prime})$. The component index selects the respective expert to obtain $\mathbf{a}^{\prime}\sim q^{\nu_{z}}(\cdot|\mathbf{s}^{\prime},z^{\prime})$. For Gaussian MoE the experts are chosen as $q^{\nu_{z}}(\mathbf{a}|\mathbf{s},z)=\mathcal{N}\left(\mathbf{a}|\mu^{\nu_{z}}(\mathbf{s}),\Sigma^{\nu_{z}}(\mathbf{s})\right)$, where $\mu^{\nu_{z}},\Sigma^{\nu_{z}}$ could be neural networks parameterized by ${\nu_{z}}$. From the properties of Gaussian distributions, it directly follows that this model class admits tractable likelihoods and fast sampling routines. Furthermore, given enough components, Gaussian MoE s are universal approximators of densities [49], which makes them a good representation for distillation of diffusion policies.
@@ 4 Variational Distillation of Denoising Diffusion Policies
with batch size $M\leq N$. Thus, $J(\phi)$ can be minimized using gradient-based optimization techniques with a gradient estimator such as reinforce [52, 53] or the reparameterization trick [54]. Note that, while the states are sampled from the given data set of the behavior policy, the actions needed to evaluate the KL are generated using our estimated model $q^{\phi}(\mathbf{a}|\mathbf{s}_{i})$. Yet, there are two difficulties to directly apply this scheme in distilling a diffusion model into an MoE: i) we are not able to evaluate the likelihood $\pi(\mathbf{a}|\mathbf{s})$ of a diffusion model, ii) training of MoE notoriously difficult [15]. We will address these two issues in Section 4.1 and Section 4.2, respectively.
@@ 4.1 Scalable Variational Inference for Denoising Diffusion Policy Distillation
Although we cannot directly evaluate the likelihood of the diffusion policy $\mathcal{\pi}(\mathbf{a}|\mathbf{s})$, we have access to its score functions $\nabla_{\mathbf{a}_{t}}\log\pi_{t}(\mathbf{a}_{t}|\mathbf{s})=\bm{f}_{{{\theta}}}(\mathbf{a}_{t},\mathbf{s},t)$, where ${t\in[0,T]}$ is the diffusion time step. In practice, we would like to evaluate the score in the limit of $t\rightarrow 0$ as $\nabla_{\mathbf{a}}\log\pi^{*}(\mathbf{a}|\mathbf{s})\approx\lim_{t\rightarrow 0}\bm{f}_{{{\theta}}}(\mathbf{a}_{t},\mathbf{s},t).$ Yet, this might lead to an unstable optimization [55] as this score is often not estimated well throughout the action space, and, hence other diffusion time-step selection processes are needed [44]. For now, we will omit the diffusion time-step for the sake of simplicity and refer to Section 4.3 for a detailed discussion about time-step selection. Moreover, we will, for now, assume that the parametrization of $q^{\phi}$ is amendable to the reparameterization trick [54], which is however not the case for mixture models. In this case, we can express $\mathbf{a}\sim q^{\phi}(\cdot|\mathbf{s})$ using a transformation $\bm{h}^{\phi}(\bm{{\epsilon}},\mathbf{s})$ with an auxiliary variable $\bm{{\epsilon}}\sim p(\cdot)$ such that $\mathbf{a}=\bm{h}^{\phi}(\bm{{\epsilon}},\mathbf{s})$. We then express the gradient of $J$ w.r.t. $\phi$ as
@@ 4.2 Variational Inference via Mixture of Experts
To distill multimodal distributions learned by diffusion models, we require a more complex family of distributions than conditional diagonal Gaussian distributions, which are commonly used in amortized VI. We will therefore use Gaussian mixture of experts. Directly computing Eq.(6) with $q^{\phi}$ being an MoE model (see Eq. 3) is difficult as mixture models do not allow a direct application of the reparametrization trick. To that end, we construct an upper bound of $J$ which is decomposible into single-objectives per component, which again allows the use of the reparametrization trick. The upper bound $U(\phi,\tilde{q})$ can be obtained by making use of the chain rule for KL divergences [56, 57, 15], i.e.,
@@
Yet, applying the reparameterization trick for $q^{\xi}$ is not straightforward. Using other gradient estimators such as reinforce [53], requires evaluating $U_{z}^{\mathbf{s}}(\nu_{z},\tilde{q})$ which is not possible as we do not have access to $\log\pi(\mathbf{a}|\mathbf{s})$. This motivates the need for a different optimization scheme. We note that $q^{\xi}$ is a categorical distribution and can approximate any distribution over $z$. It does therefore not suffer from problems associated with the M-projection such as mode averaging due to limited complexity of $q^{\xi}$. We thus propose to use the forward KL divergence for optimizing $\xi$, i.e.,
@@
and, therefore does not affect the convergence guarantees of our method.
@@ 5 Experiments
DDPM BESO DDPM-1 BESO-1 CD VDD-DDPM(ours) VDD-BESO(ours)
Kitchen - $4.06$ $\underline{4.02}$ $\bm{3.87\pm 0.05}$ $3.85\pm 0.10$
Block Push - $0.96$ $\underline{0.94}$ $0.89\pm 0.05$ $\underline{\bm{0.91\pm 0.03}}$
Avoiding $0.94$ $0.96$ $0.09$ $0.84$ $0.82\pm 0.05$ $0.92\pm 0.02$ $\bm{\underline{0.95\pm 0.01}}$
Aligning $0.85$ $0.85$ $0.00$ $0.93$ $\bm{\underline{0.94\pm 0.08}}$ $0.70\pm 0.07$ $0.86\pm 0.04$
Pushing $0.74$ $0.78$ $0.00$ $0.70$ $0.66\pm 0.05$ $0.61\pm 0.04$ $\underline{\bm{0.85\pm 0.02}}$
Stacking-1 $0.89$ $0.91$ $0.00$ $0.75$ $0.69\pm 0.06$ $0.81\pm 0.08$ $\underline{\bm{0.85\pm 0.02}}$
Stacking-2 $0.68$ $0.70$ $0.00$ $0.53$ $0.46\pm 0.11$ $\underline{\bm{0.60\pm 0.07}}$ $0.57\pm 0.06$
Sorting (Image) $0.69$ $0.70$ $0.20$ $0.68$ $0.71\pm 0.07$ $\underline{\bm{0.80\pm 0.04}}$ $0.76\pm 0.04$
Stacking (Image) $0.58$ $0.66$ $0.00$ $0.58$ $0.63\pm 0.01$ $\underline{\bm{0.78\pm 0.02}}$ $0.60\pm 0.04$
@@
DDPM BESO DDPM-1 BESO-1 CD VDD-DDPM(ours) VDD-BESO(ours)
Avoiding $0.89$ $0.87$ $0.25$ $\underline{0.76}$ $\bm{0.72\pm 0.02}$ $0.37\pm 0.01$ $\bm{0.72\pm 0.12}$
Aligning $0.62$ $0.67$ $0.00$ $0.34$ $0.32\pm 0.14$ $0.25\pm 0.09$ $0.40\pm 0.04$
Pushing $0.74$ $0.76$ $0.00$ $0.50$ $0.53\pm 0.07$ $0.66\pm 0.05$ $\bm{0.69\pm 0.08}$
Stacking-1 $0.24$ $0.30$ $0.00$ $0.26$ $\bm{0.19\pm 0.12}$ $\bm{0.19\pm 0.05}$ $0.16\pm 0.03$
Stacking-2 $0.12$ $0.13$ $0.00$ $0.07$ $0.03\pm 0.05$ $0.07\pm 0.04$ $\bm{0.13\pm 0.06}$
Sorting (Image) $0.16$ $0.19$ $0.09$ $0.14$ $0.14\pm 0.06$ $0.12\pm 0.03$ $\bm{0.22\pm 0.03}$
Stacking (Image) $0.31$ $0.15$ $0.00$ $0.10$ $0.06\pm 0.01$ $0.05\pm 0.02$ $\bm{0.11\pm 0.03}$
@@
We conducted imitation learning experiments by distilling two types of diffusion models: variance preserving (VP) [2, 12] and variance exploding (VE) [59, 4]. We selected DDPM as the representative for VP and BESO as the representative for VE. In the experiments, DDPM-1 and BESO-1 denote the results when performing only one denoising step of the respective diffusion models during inference. VDD-DDPM and VDD-BESO denote the results of distilled VDD Additionally, we consider the SoTA Consistency Distillation (CD) [40] as a baseline for comparing VDD’s performance in distillation. We compare VDD against MoE learning baselines, namely the widely-used Expectation-Maximization (EM) [60] approach as a representative of the maximum likelihood-based objective and the recently introduced SoTA method Information Maximizing Curriculum (IMC) as representative of the reverse KL-based objective. To make them stronger baselines, we extend them with the architecture described in Figure 5 and name the extended methods as EM-GPT and IMC-GPT, respectively. For a fair comparison, we used the same diffusion models as the origin model for all distillation methods that we have trained on seed 0. For a statistically significant comparison, all methods have been run on 4 random seeds, and the mean and the standard deviation are reported throughout the evaluation.
The evaluations are structured as follows. Firstly, we demonstrate that VDD is able to achieve competitive performance with the SoTA diffusion distillation method and the original diffusion models on two established datasets. Next, we proceed to a recently proposed challenging benchmark with human demonstrations, where VDD outperforms existing diffusion distillation and SoTA MoE learning approaches. Following this, we highlight the faster inference time of VDD. Finally, a series of ablation studies reflect the importance of VDD’s essential algorithmic properties.

### v2

@@ Abstract
This work introduces Variational Diffusion Distillation (VDD), a novel method that distills denoising diffusion policies into Mixtures of Experts (MoE) through variational inference. Diffusion Models are the current state-of-the-art in generative modeling due to their exceptional ability to accurately learn and represent complex, multi-modal distributions. This ability allows Diffusion Models to replicate the inherent diversity in human behavior, making them the preferred models in behavior learning such as Learning from Human Demonstrations (LfD). However, diffusion models come with some drawbacks, including the intractability of likelihoods and long inference times due to their iterative sampling process. The inference times, in particular, pose a significant challenge to real-time applications such as robot control. In contrast, MoEs effectively address the aforementioned issues while retaining the ability to represent complex distributions but are notoriously difficult to train. VDD is the first method that distills pre-trained diffusion models into MoE models, and hence, combines the expressiveness of Diffusion Models with the benefits of Mixture Models. Specifically, VDD leverages a decompositional upper bound of the variational objective that allows the training of each expert separately, resulting in a robust optimization scheme for MoEs. VDD demonstrates across nine complex behavior learning tasks, that it is able to: i) accurately distill complex distributions learned by the diffusion model, ii) outperform existing state-of-the-art distillation methods, and iii) surpass conventional methods for training MoE. The code and videos are available at https://intuitive-robots.github.io/vdd-website.
@@ 1 Introduction
To obtain the benefits of both models, i.e., learning highly accurate generative models with diffusion and obtaining simple, tractable models using a mixture of experts, this work introduces Variational Diffusion Distillation (VDD), a novel method that distills diffusion models to MoEs. Starting from the variational inference objective [17, 18], we derive a lower bound that decomposes the objective into separate per-expert objectives, resulting in a robust optimization scheme. Each per-expert objective elegantly leverages the gradient of the pre-trained score function such that the MoE benefits from the diffusion model’s properties. The resulting MoE policy performs on par with the diffusion model and covers the same modes, while being interpretable, faster during inference, and has a tractable likelihood. This final policy is readily available to the user for post hoc analysis or fast fine-tuning for more specific situations. A high-level architecture of the VDD model and its relation to the diffusion policy is shown in Fig. 1. VDD is thoroughly evaluated on nine complex behavior-learning tasks that demonstrate the aforementioned properties. As an additional insight, this paper observed that one-step continuous diffusion models already perform well, a finding not discussed in prior work.
@@ 2 Related Works
Knowledge Distillation from Diffusion Models Knowledge distillation from diffusion models has been researched in various research areas. For instance, in text-to-3D modeling, training a NeRF-based text-to-3D model without any 3D data by mapping the 3D scene to a 2D image and leveraging a text-to-2D diffusion is proposed [7]. The work proposes minimizing the Score Distillation Sampling (SDS) loss that is inspired by probability density distillation [40] and incentivizes the 3D-model to be updated into higher density regions as indicated by the score function of the diffusion model. To overcome drawbacks such as over-smoothing and low-diversity problems when using the SDS loss, variational score distillation (VSD) treats the 3D scene as a random variable and optimizes a distribution over these scenes such that the projected 2D image aligns with the 2D diffusion model [8]. In a similar context, the work in [41] proposes distilling a trained diffusion model into another diffusion model while progressively reducing the number of steps. However, even though the number of diffusion steps is drastically reduced, a complete distillation, i.e. one-step inference as for VDD is not provided. Additionally, the resulting model suffers from the same drawbacks of diffusion models such as intractable likelihoods. In contrast, in consistency distillation (CD), diffusion models are distilled to consistency models (CM) [42, 43, 44, 45] such that data generation is possible in one step from noise to data. However, one-step data generation typically results in lower sample quality, requiring a trade-off between iterative and single-step generation based on the desired outcome. As CMs, VDD performs one-step data generation but distills the pre-trained diffusion model to an MoE which has a tractable likelihood and is efficient in inference time. The experimental evaluations show the advantages of VDD over CMs. Diff-Instruct [46] proposes a two-step framework for distilling diffusion models into implicit generative models, whereas VDD considers an explicit generative model where the model’s density can be directly evaluated. In addition, Diff-Instruct requires training an auxiliary diffusion model, while VDD only optimizes a single model. Score Regularized Policy Optimization (SRPO) [47] also leverages a diffusion behavior policy to regularize the offline RL-based objective. However, in contrast to SRPO, VDD learns an MoE policy instead of a uni-modal Gaussian policy and explicitly distills a diffusion model instead of using it as guidance during optimization. Furthermore, VDD trains MoEs policies in imitation learning instead of reward-labeled data as in offline RL. A concurrent work, EM-Distillation (EMD)[48], introduces an EM-style distillation objective derived from the mode-covering forward KL divergence. In contrast, VDD proposes an EM-style objective based on the mode-seeking reverse KL but encourages mode-covering behavior by having multiple experts.
@@ 3 Preliminaries
where $q^{\xi}(z|\mathbf{s})$ and $q^{\nu_{z}}(\mathbf{a}|\mathbf{s},z)$ are referred to as gating and experts respectively. $\xi$ and $\nu_{z}$ denote the gating and expert parameters and thus $\phi=\xi\cup\{\nu_{z}\}_{z}$. The gating is responsible for soft-partitioning the state space into sub-regions where the corresponding experts approximate the target density. To sample actions, that is, $\mathbf{a}^{\prime}\sim q^{\phi}(\cdot|\mathbf{s}^{\prime})$ for some state $\mathbf{s}^{\prime}$, we first sample a component index from the gating, i.e., $z^{\prime}\sim q^{\xi}(\cdot|\mathbf{s}^{\prime})$. The component index selects the respective expert to obtain $\mathbf{a}^{\prime}\sim q^{\nu_{z}}(\cdot|\mathbf{s}^{\prime},z^{\prime})$. For Gaussian MoE the experts are chosen as $q^{\nu_{z}}(\mathbf{a}|\mathbf{s},z)=\mathcal{N}\left(\mathbf{a}|\mu^{\nu_{z}}(\mathbf{s}),\Sigma^{\nu_{z}}(\mathbf{s})\right)$, where $\mu^{\nu_{z}},\Sigma^{\nu_{z}}$ could be neural networks parameterized by ${\nu_{z}}$. From the properties of Gaussian distributions, it directly follows that this model class admits tractable likelihoods and fast sampling routines. Furthermore, given enough components, Gaussian MoE s are universal approximators of densities [53], which makes them a good representation for distillation of diffusion policies.
@@ 4 Variational Distillation of Denoising Diffusion Policies
with batch size $M\leq N$. Thus, $J(\phi)$ can be minimized using gradient-based optimization techniques with a gradient estimator such as reinforce [56, 57] or the reparameterization trick [58]. Note that, while the states are sampled from the given data set of the behavior policy, the actions needed to evaluate the KL are generated using our estimated model $q^{\phi}(\mathbf{a}|\mathbf{s}_{i})$. Yet, there are two difficulties to directly apply this scheme in distilling a diffusion model into an MoE: i) we are not able to evaluate the likelihood $\pi(\mathbf{a}|\mathbf{s})$ of a diffusion model, ii) training of MoE models is notoriously difficult [15]. We will address these two issues in Section 4.1 and Section 4.2, respectively.
@@ 4.1 Scalable Variational Inference for Denoising Diffusion Policy Distillation
Although we cannot directly evaluate the likelihood of the diffusion policy $\mathcal{\pi}(\mathbf{a}|\mathbf{s})$, we have access to its score functions $\nabla_{\mathbf{a}_{t}}\log\pi_{t}(\mathbf{a}_{t}|\mathbf{s})=\bm{f}_{{{\theta}}}(\mathbf{a}_{t},\mathbf{s},t)$, where ${t\in[0,T]}$ is the diffusion time step. In practice, we would like to evaluate the score in the limit of $t\rightarrow 0$ as $\nabla_{\mathbf{a}}\log\pi^{*}(\mathbf{a}|\mathbf{s})\approx\lim_{t\rightarrow 0}\bm{f}_{{{\theta}}}(\mathbf{a}_{t},\mathbf{s},t).$ Yet, this might lead to an unstable optimization [59] as this score is often not estimated well throughout the action space, and, hence other diffusion time-step selection processes are needed [47]. For now, we will omit the diffusion time-step for the sake of simplicity and refer to Section 4.3 for a detailed discussion about time-step selection. Moreover, we will, for now, assume that the parametrization of $q^{\phi}$ is amendable to the reparameterization trick [58]. In this case, we can express $\mathbf{a}\sim q^{\phi}(\cdot|\mathbf{s})$ using a transformation $\bm{h}^{\phi}(\bm{{\epsilon}},\mathbf{s})$ with an auxiliary variable $\bm{{\epsilon}}\sim p(\cdot)$ such that $\mathbf{a}=\bm{h}^{\phi}(\bm{{\epsilon}},\mathbf{s})$. We then express the gradient of $J$ w.r.t. $\phi$ as
@@ 4.2 Variational Inference via Mixture of Experts
To distill multimodal distributions learned by diffusion models, we require a more complex family of distributions than conditional diagonal Gaussian distributions, which are commonly used in amortized VI. We will therefore use Gaussian mixture of experts. To that end, we construct an upper bound of $J$ which is decomposable into single objectives per expert, allowing for reparameterizing each expert individually and therefore avoiding the need for techniques that perform reparameterization for the entire MoE [60, 61]. The upper bound $U(\phi,\tilde{q})$ can be obtained by making use of the chain rule for KL divergences [62, 63, 15], i.e.,
@@
Using gradient estimators such as reinforce [57], requires evaluating $U_{z}^{\mathbf{s}}(\nu_{z},\tilde{q})$ which is not possible as we do not have access to $\log\pi(\mathbf{a}|\mathbf{s})$. This motivates the need for a different optimization scheme. We note that $q^{\xi}$ is a categorical distribution and can approximate any distribution over $z$. It does therefore not suffer from problems associated with the forward KL divergence such as mode averaging due to limited complexity of $q^{\xi}$. We thus propose using the following objective for optimizing $\xi$, i.e.,
@@
and, therefore does not affect the convergence guarantees of our method. Please note that the forward KL requires samples from the teacher model $\pi$. In practice, if the dataset used for training the diffusion model is available, it can be used as a proxy and prevent costly data regeneration.
@@ 5 Experiments
VP (DDPM) VE (BESO) VP-1 VE-1 CD-VE CTM-VE VDD-VP(ours) VDD-VE(ours)
Kitchen $3.35$ $4.06$ $0.22$ $\underline{4.02}$ $3.87\pm 0.05$ $\bm{3.89\pm 0.11}$ $3.24\pm 0.12$ $3.85\pm 0.10$
Block Push $0.96$ $0.96$ $0.09$ $\underline{0.94}$ $0.89\pm 0.05$ $0.89\pm 0.04$ $0.93\pm 0.03$ $\underline{\bm{0.91\pm 0.03}}$
Avoiding $0.94$ $0.96$ $0.09$ $0.84$ $0.82\pm 0.05$ $0.93\pm 0.02$ $0.92\pm 0.02$ $\bm{\underline{0.95\pm 0.01}}$
Aligning $0.85$ $0.85$ $0.00$ $0.93$ $\bm{\underline{0.94\pm 0.08}}$ $0.81\pm 0.11$ $0.70\pm 0.07$ $0.86\pm 0.04$
Pushing $0.74$ $0.78$ $0.00$ $0.70$ $0.66\pm 0.05$ $0.80\pm 0.07$ $0.61\pm 0.04$ $\underline{\bm{0.85\pm 0.02}}$
Stacking-1 $0.89$ $0.91$ $0.00$ $0.75$ $0.69\pm 0.06$ $0.54\pm 0.17$ $0.81\pm 0.08$ $\underline{\bm{0.85\pm 0.02}}$
Stacking-2 $0.68$ $0.70$ $0.00$ $0.53$ $0.46\pm 0.11$ $0.30\pm 0.09$ $\underline{\bm{0.60\pm 0.07}}$ $0.57\pm 0.06$
Sorting (Image) $0.69$ $0.70$ $0.20$ $0.68$ $0.71\pm 0.07$ $0.70\pm 0.07$ $\underline{\bm{0.80\pm 0.04}}$ $0.76\pm 0.04$
Stacking (Image) $0.58$ $0.66$ $0.00$ $0.58$ $0.63\pm 0.01$ $0.59\pm 0.10$ $\underline{\bm{0.78\pm 0.02}}$ $0.60\pm 0.04$
@@
VP (DDPM) VE (BESO) VP-1 VE-1 CD-VE CTM-VE VDD-VP(ours) VDD-VE(ours)
Avoiding $0.89$ $0.87$ $0.25$ $0.76$ $0.72\pm 0.02$ $\bm{0.79\pm 0.04}$ $0.37\pm 0.01$ $0.72\pm 0.12$
Aligning $0.62$ $0.67$ $0.00$ $0.34$ $0.32\pm 0.14$ $0.31\pm 0.28$ $0.25\pm 0.09$ $0.40\pm 0.04$
Pushing $0.74$ $0.76$ $0.00$ $0.50$ $0.53\pm 0.07$ $0.54\pm 0.08$ $0.66\pm 0.05$ $\bm{0.69\pm 0.08}$
Stacking-1 $0.24$ $0.30$ $0.00$ $0.26$ $\bm{0.19\pm 0.12}$ $0.18\pm 0.08$ $\bm{0.19\pm 0.05}$ $0.16\pm 0.03$
Stacking-2 $0.12$ $0.13$ $0.00$ $0.07$ $0.03\pm 0.05$ $0.09\pm 0.06$ $0.07\pm 0.04$ $\bm{0.13\pm 0.06}$
Sorting (Image) $0.16$ $0.19$ $0.09$ $0.14$ $0.14\pm 0.06$ $0.08\pm 0.05$ $0.12\pm 0.03$ $\bm{0.22\pm 0.03}$
Stacking (Image) $0.31$ $0.15$ $0.00$ $0.10$ $0.06\pm 0.01$ $0.04\pm 0.04$ $0.05\pm 0.02$ $\bm{0.11\pm 0.03}$
@@
We conducted imitation learning experiments by distilling two types of diffusion models: variance preserving (VP) [2, 12] and variance exploding (VE) [65, 4]. We selected DDPM as the representative for VP and BESO as the representative for VE. In the experiments, DDPM-1 and BESO-1 denote the results when performing only one denoising step of the respective diffusion models during inference. VDD-DDPM and VDD-BESO denote the results of distilled VDD Additionally, we consider the SoTA Consistency Distillation (CD) [42] and Consistency Trajectory Model (CTM) [44] as baselines for comparing VDD’s performance in distillation. For CD and CTM, we distill from the VE following the original works. For CTM we adapt the implementation and design choices from Consistency Policy [45], which are specialized for behavior learning. We compare VDD against MoE learning baselines, namely the widely-used Expectation-Maximization (EM) [66] approach as a representative of the maximum likelihood-based objective and the recently introduced SoTA method Information Maximizing Curriculum (IMC) as representative of the reverse KL-based objective. To make them stronger baselines, we extend them with the architecture described in Figure 6 and name the extended methods as EM-GPT and IMC-GPT, respectively. For a fair comparison, we used the same diffusion models as the origin model for all distillation methods that we have trained on seed 0. For a statistically significant comparison, all methods have been run on 4 random seeds, and the mean and the standard deviation are reported throughout the evaluation. Detailed descriptions regarding the baselines implementation and hyperparameters selection can be found in Appendix D and E.
The evaluations are structured as follows. Firstly, we demonstrate that VDD is able to achieve competitive performance with the SoTA diffusion distillation method and the original diffusion models on two established datasets. Next, we proceed to a recently proposed challenging benchmark with human demonstrations, where VDD outperforms existing diffusion distillation and SoTA MoE learning approaches. We then highlight the faster inference time of VDD. Following this, a series of ablation studies reflect the importance of VDD’s essential algorithmic properties. Finally, we provide a visualization to offer deeper insights into our method.